OpenStudy (anonymous):

Integral sin^3x-cos^2x/sin^2x dx

OpenStudy (anonymous):

i got this: sinx dx-cot

OpenStudy (anonymous):

is sin^2(x) dividing the whole thing?

OpenStudy (anonymous):

yes sin^2x is the denominator

OpenStudy (bahrom7893):

Simplify and integrate..

OpenStudy (anonymous):

\[((\sin^3(x)-\cos^2(x))/\sin^2(x))dx ?\]